1/2x + 1/2y = 2
1/3x - 1/3y = 4

Answer by Alan3354(69443)

1/2x + 1/2y = 2
1/3x - 1/3y = 4
-----------------
Assuming 1/2x means (1/2)x and not 1/(2x):
Multiply the 1st eqn by 2 and the 2nd by 3
x + y = 4
x - y = 12
----------- Add
2x = 16
you can finish it.
---------------
Use parentheses to eliminate the ambiguities.